Force is proportional to AB,BC and 2CA acts along the side of triangle ABC in magnitude and direction will be given by what?

Grade:11

1 Answers

Arun
5 years ago
Dear Roshan
 
The force on side AB = AB

Force on BC = BC

Force on AC = 2CA

1.) Magnitude of AB = AC + CB

= 2CA - BC

2.) magnitude of BC = BA + AC

= 2CA - AB

3.) magnitude of AC = AB + BC

= AB + BC